# EpochCore: Digital Hardware Accelerator For Structured State-Space Models

### v1

@@ IV-C1 Energy and Latency Comparison
Figure 11: Energy/Inference for executing Liquid-S4 and DNN layers for various LRA datasets shown in Table I when using TPU + EpochCore and only EpochCore accelerators.
Despite higher energy costs for DNN layers, EpochCore’s efficiency and throughput gains in S4 layer inference (Figure 10) make it a better choice for real-world applications.
@@
IV-D Full model evaluation of EpochCore
In Figure 15 the latency breakdown between S4 and DNN/SNN layers (as in Figure 3) are shown for evaluations on GPU (Nvidia A100), Sparse-SA and EpochCore accelerators for the same workload. The EpochCore accelerator demonstrates significant improvement in S4 layer latency compared to Sparse-SA. The breakdown also reveals that EpochCore significantly reduces the overhead of S4 layers from $\sim$95to $\sim$7%, allowing for a more balanced and efficient execution across different layers. Similarly, latency breakdown on additional LRA datasets using the time-variant Liquid-S4 model is shown in Figure 13. The average latency gain for inference of Liquid-S4 layer is shown to be around 1,666×1,666$\times$1 , 666 × to 16,270×16,270$\times$16 , 270 × as shown in Figure 14.
Figure 15: Latency breakdown of S4-based models for CIFAR-10 dataset with sequence length, T=64K, using GPU, Sparse-SA and EpochCore accelerators.
TABLE V: DNN Experiments
Experiment
@@ IV-D Full model evaluation of EpochCore
IV-E Ablation Study
@@
IV-E1 Impact of S4 state-map size on accuracy
In Figure 16 the accuracy impact after 10-training iterations is show for various datasets while varying the hidden state-map size. PathFinger testcase require much large training iterations to show the sensitivity of state-map size. For other datasets, choosing the state-map size impacts accuracy. Most SoTA model evaluations pick the state-map size of $N=64$, though some datasets may require higher dimensionality.
Figure 16: Accuracy of S4 model training with different hidden state-map size $N$.
IV-E2 On-Chip SRAM Bandwidth Comparison
@@
The maximum on-chip SRAM size for S4-based workloads is determined by the input data sequence length and input batch size. As shown in Figure 18, an input sequence of 1-million, typical for audio and other LRA datasets, can fit within 10MB SRAM with 1-batch input, in a $\times 64$ 64 SA with 32-bit data. Shorter sequence lengths allow for larger batch sizes within the same SRAM capacity. The key observations are: (1) Input/Output SRAMs dominate overall memory usage, and are 2-3 orders of magnitude larger than Weight SRAM; (2) the required Weight SRAM size does not depend on the sequence length; and (3) given a maximum of 10MB SRAM, the figure shows the feasible combinations of batch sizes and sequence lengths (for example a batch size of 32 can support sequence length upto 64K) within a single SSM layer that can be supported at a time by EpochCore.
@@
IV-E3 PE Utilization
@@ V Related Work
EpochCore is the first ever unified accelerator to support multiple structured-SSMs as well as traditional GEMM workloads. Recent advances in structured SSMs have led to a surge of hardware accelerators (MARCA [19], VGA [17], FastMamba [34]) tailored to support long-context sequence modeling. These accelerators vary in their choice of computational paradigm—recurrence vs. convolution—and in their specialization toward specific SSM variants. We categorize and compare these efforts in relation to our proposed design, EpochCore.
@@
EpochCore offers a more targeted and efficient architecture for accelerating structured SSMs, by directly exploiting the recurrent nature of SSM computations rather than relying on convolution-based computations. While MARCA [21] uses a unified PE array and buffers to support both linear and element-wise operations in Mamba models, EpochCore introduces specialized compute units (e.g., LIMA-PE) and programmable dataflows (ProDF) tailored for real-time streaming and low-bitwidth recurrence operations, enabling significantly higher performance and energy efficiency. EpochCore is more generalizable than VGA [17] as extensions to support Mamba [7] are possible through incremental updates to LIMA-PE to support new MAC operations for gated layer and improvements to ProDF to support additional layers for variable discretization. The deep integration of SSM-specific primitives—such as exponential decay, real/complex MACs, and minimal memory movement—allows EpochCore to achieve over 2000×2000$\times$2000 × speedups compared to GPUs on S4-based CIFAR-10 dataset with sequence length of $64K$, while also offering extensibility to support a broader class of SSM models beyond Mamba.
@@ VI Conclusion
In this paper we present EpochCore, a novel digital hardware accelerator, to infer S4 and Liquid-S4 based neural models besides general GEMM-based networks. We proposed, designed, and evaluated a novel processing element, LIMA-PE. We proposed a novel dataflow, ProDF, to programmatically configure various dataflow tailored for S4, Liquid-S4 and GEMM dataflows. The performance and energy efficiency of EpochCore was evaluated against other accelerators. The EpochCore delivers 250×250$\times$250 × performance improvement and 45×45$\times$45 × reduction in total energy consumption compared to traditional PE-based Sparse SA accelerators for Liquid-S4 based models and around 2,000$\sim$ 2,000$\times$ 2 , 000 × improvement in latency/inference on LRA datasets compared to GPU kernel operations. Additionally, EpochCore achieves 25×25$\times$25 × speed-up and 10×10$\times$10 × reduction in total energy consumption compared to Boriakoff’s 1-D SA across the same LRA datasets. EpochCore also reduces memory bandwidth usage by 3×3$\times$3 × compared to Boriakoff’s 1-D SA. The LIMA-PE design incurs a 1.42×1.4-2$\times$1.4 - 2 × increase in area and up to 1.6×1.6$\times$1.6 × increase in power compared to Traditional-PEs. However, due to the use of a gated-clock PE design, power savings of 1.63×1.6-3$\times$1.6 - 3 × are observed between different operational modes. A comparative latency breakdown analysis between S4 and DNN/CNN layers shows that EpochCore scales well with sequence length maintaining an even distribution of workload across layers, unlike GPU-based implementations where S4 layers tend to dominate.

### v2

@@ IV-C1 Energy and Latency Comparison
Figure 11: Energy/Inference for executing Liquid-S4 and DNN layers for various LRA datasets shown in Table I when using TPU + EpochCore and only EpochCore accelerators. The overhead of EpochCore on DNN layers is negligible.
Despite higher energy costs for DNN layers, EpochCore ’s efficiency and throughput gains in Liquid-S4 layer inference (Figure 10) make it a better choice for real-world applications.
@@
IV-D Full model evaluation of EpochCore
In Figure 13 the latency breakdown between S4 and DNN/SNN layers (as in Figure 3) are shown for evaluations on GPU (Nvidia A100), Sparse-SA and EpochCore accelerators for the same workload. The EpochCore accelerator demonstrates significant improvement in S4 layer latency compared to Sparse-SA. The breakdown also reveals that EpochCore significantly reduces the latency of S4 layers from $\sim$95to $\sim$7%, allowing for a more balanced and efficient execution across different layers. Similarly, latency breakdown on additional LRA datasets using the time-variant Liquid-S4 model is shown in Figure 14 with the corresponding Liquid-S4 model parameters shown in Table I. The average latency gain for inference of Liquid-S4 layer is shown to be around 1,666×1,666$\times$1 , 666 × to 16,270×16,270$\times$16 , 270 × as shown in Figure 15.
Figure 13: Latency breakdown of S4 based models for CIFAR-10 dataset with sequence length, T=64K, using GPU, Sparse-SA and EpochCore accelerators.
@@ IV-D Full model evaluation of EpochCore
TABLE V: DNN Data Sets and Model Parameters.
Dataset
@@
IV-E EpochCore vs Other SSM Accelerators
Table  VI shows up to 3860$\times$ speedup for H3-SSM models on EpochCore, outperforming VGA [17]. For Mamba models, EpochCore is closely comparable to other accelerators such as FastMamba [34] and MARCA [19]. EpochCore is the first accelerator for Liquid-S4 models.
IV-F Ablation Study
@@
IV-F1 Impact of Liquid-S4 state-map size on accuracy
In Figure 16 the accuracy impact after ten training iterations is shown for various datasets while varying the hidden state-map size. PathFinder requires much larger training iterations to show the sensitivity of state-map size. The state map size has no impact on the accuracy of Pathfinder. For other datasets, choosing the state-map size impacts accuracy. Most SoTA models pick the state-map size of $N=64$.
Figure 16: Accuracy of Liquid-S4 model training with different hidden state-map size $N$.
IV-F2 On-Chip SRAM Bandwidth and Size Comparison
@@
The maximum on-chip SRAM size for S4, Liquid-S4 based workloads is determined by the input data sequence length and input batch size. As shown in Figure 18, an input sequence of 1-million, typical for audio and other LRA datasets, can fit within 10MB SRAM with 1-batch input, in a $\times 64$ 64 SA with 32-bit data. Shorter sequence lengths allow for larger batch sizes within the same SRAM capacity. The key observations are: (1) Input/Output SRAMs dominate overall memory usage, and are 2-3 orders of magnitude larger than Weight SRAM; (2) the required Weight SRAM size does not depend on the sequence length; and (3) given a maximum of 10MB SRAM, the figure shows the feasible combinations of batch sizes and sequence lengths (for example a batch size of 32 can support sequence length upto 64K) within a single SSM layer that can be supported at a time by EpochCore.
@@
TABLE VI: Performance Speed up over GPU
SSM-Accelerator H3-SSM Mamba-SSM Liquid-S4
VGA[17] 4$\times$ - -
MARCA[19] - 11.66$\times$ -
FastMamba[34] - 6.06$\times$ -
EpochCore 3860$\times$ 4.75$\times$ 2000$\times$
IV-F3 PE Utilization
@@ V Related Work
TABLE VII: Comparison of SSM accelerators
Metric FastMamba VGA Marca EpochCore
SSM Mamba2 H3 Mamba S4/Liquid-S4
Time-Variant Yes No Yes Yes
Complex value No Yes No Yes
Speed up over GPU 8.9X 14.9X 11.6X 2000X
Recent advances in structured SSMs have led to a surge of hardware accelerators (MARCA [19], VGA [17], FastMamba [34]) tailored to support long-context sequence modeling. These accelerators vary in their choice of computational paradigm—recurrence vs. convolution—and in their specialization toward specific SSM variants. We categorize and compare these efforts in relation to our EpochCore.
@@
EpochCore is the first unified accelerator supporting multiple structured SSMs (S4, Liquid-S4, H3, Mamba) and GEMM-based DNN layers. Unlike convolution-based designs, EpochCore exploits the recurrent structure of SSMs via novel LIMA-PE modes and a programmable dataflow (ProDF). It supports real and complex MACs, exponential decay primitives, and weight-stationary data reuse, enabling 2000$\times$ speedup over GPUs for S4 and Liquid-S4 models.
Compared to VGA and MARCA, EpochCore is more extensible: gated Mamba variants can be supported through incremental LIMA-PE and ProDF enhancements. Its generality and high efficiency make it well-suited for diverse SSM and hybrid deep model workloads.
@@ VI Conclusion
This paper presents EpochCore, a novel digital accelerator for S4 and Liquid-S4 inference, as well as general DNN workloads. We introduce two key innovations: the LIMA-PE, a specialized processing element supporting real and complex recurrent operations, and ProDF, a programmable dataflow optimized for SSMs and GEMM layers.
EpochCore achieves up to 250$\times$ speedup and 45$\times$ energy savings over Sparse SA accelerators for Liquid-S4 models, and 25$\times$ faster and 10$\times$ more energy-efficient than Boriakoff’s 1D SA. Compared to GPUs, EpochCore delivers  2000$\times$ performance gains on LRA datasets using S4 and Liquid-S4 models, while reducing memory bandwidth usage by 3$\times$. On other structured SSM models such as H3 and Mamba, EpochCore shows performance gains of 3860$\times$ and 4.75$\times$ respectively over GPUs.
LIMA-PE incurs a 1.4–2$\times$ area and up to 1.6$\times$ power overhead versus traditional PEs, but gated-clock design enables 1.6–3$\times$ power savings across operating modes. Latency analysis further shows that EpochCore maintains balanced execution between S4/Liquid-S4 and DNN layers, scaling efficiently with sequence length—unlike GPU-based systems where S4/Liquid-S4 dominates runtime.